# ED-ViT: Splitting Vision Transformer for Distributed Inference on Edge Devices

### v1

@@ 4. Methodology
This section describes the design of the ED-ViT framework proposed to solve the optimization problem outlined in Eq. 1. We first explain the main workflow of ED-ViT and then provide detailed descriptions of the four key steps involved.
4.1. Design Overview
@@
As illustrated in Figure 1, ED-ViT leverages the unique characteristics of Vision Transformer and the collaboration of multiple edge devices. The framework involves $N$ concurrent edge devices for distributed inference alongside a lightweight MLP aggregation to derive the final classification results. Initially, the original Vision Transformer is trained on the entire dataset to achieve high test accuracy for the classification task. The ED-ViT framework is composed of four main components: model splitting, pruning, assignment, and fusion. During model splitting, the Vision Transformer model is divided into sub-models, each responsible for a subset of classes. To reduce computation overhead, these sub-models are further pruned using model pruning techniques. Subsequently, the sub-models are assigned to the appropriate edge devices, taking the optimization problem into consideration. Finally, the aggregation server fuses the outputs from the edge devices to produce the final inference results. The specific details of each component are provided below.
@@
5:  until $\forall C_{a},C_{b}\in C,\big{|}|C_{a}|-|C_{b}|\big{|}\leq 1$.
6:  while $flag_{r}$ is $True$ do
7:     for $i$ in $N$ do
@@
9:     end for
10:     $MA=\phi$
11:     if $\sum_{j}m_{i}\leq bu$ then
12:        $MA=greedySearchAssign(E,M,\text{Model}_{i},D)$.
13:     end if
14:     if  $MA\neq\phi$  then
15:        $flag_{r}=False$.
16:     else
17:        $hp_{j}=hp_{j}+1$ where Modelj has the biggest memory size.
18:     end if
19:  end while
20:  $concat_{out}$ = $concat$(Model1($\mathbf{X}$),…, ModelN($\mathbf{X}$)).
21:  $MLP$ = $train$($concat_{out}$, $\mathbf{y}$).
22:  return $MLP$, Model1, …, ModelN.
4.2. Model Splitting
In the original Vision Transformer, different heads contribute to learning and inferring from the samples. However, for certain classes, maintaining all the connections between the heads can be redundant. As a result, ED-ViT prunes these connections and reconstructs the heads, with more retained heads leading to more parameters and connections being preserved. As illustrated in Algorithm 1, each Vision Transformer sub-model undergoes pruning based on a head number threshold and its associated categories, following a relatively equitable workload distribution. Subsequently, a greedy search mechanism is used to identify the most suitable edge device model assignment plan for deploying a particular sub-model, considering both energy and memory constraints. If the total memory size exceeds the budget or no suitable plan is found, an iterative approach is applied to adjust the number of heads for the sub-model with the biggest memory size to be pruned, repeating the allocation process until all sub-models are successfully assigned to edge devices. The pruning and the greedy assignment methods are shown as Algorithm 2 and Algorithm 3, located in Section 4.3 and Section 4.4, respectively.
4.3. Model Pruning
@@
We believe that reducing the computational burden of Vision Transformer will significantly contribute to lowering inference latency in distributed edge device settings. We focus on the original ViT architecture (Alexey, 2020), chosen for its simplicity and well-defined design space, focusing on redistributing the dimensionality across different blocks to achieve a more balanced tradeoff between computational efficiency and accuracy, as shown in Figure 2.
Analysis of Prunable Parameters: The main prunable components in a ViT block, as illustrated in Figure 2.
@@
Pruning Process: As illustrated in Figure 2, The pruning process is carried out in stages, with each stage focusing on one of the prunable components. We compute the KL-Divergence between the output distributions of the original model and the pruned model to evaluate the importance of each component, as follows:
@@
We focus on pruning the channels of the residual connections (shown in red) in the first stage. Using KL-Divergence, we identify and prune the channels that contribute the least, reducing the dimensionality from $d$ to $s\times d$, and the pruning factor $s$ controls the degree of reduction in the parameters. We use $j$-th sub-model as an example: we set $s=(h-hp_{j})/h$, effectively controlling the extent of the pruning and parameter reduction. This helps to streamline the flow of information between layers without significantly affecting model performance. Then, instead of directly removing entire heads in the MHSA module, we prune the least important dimensions within the query, key, and value projections ($d_{q}$, $d_{k}$, and $d_{v}$) across multiple heads. This process effectively reduces the total number of heads to $s\times h$, without entirely discarding any head, thus maintaining a balanced representation of the attention mechanism while reducing its complexity. The dimensionality of the projections is scaled accordingly to reflect the merging and pruning process, ensuring that the model retains its ability to capture token interactions. The final stage involves pruning the hidden dimension $c$ in the FFN, as shown in blue. By calculating KL-Divergence, we identify the least important neurons and reduce the hidden dimension from $c$ to $s\times c$. Following each pruning stage, the model is fine-tuned to recover any performance loss that may result from the parameter reduction. This ensures that the pruned model achieves a similar level of accuracy as the original model while requiring fewer computational resources. The details are shown in Appendix B.
@@ 4.4. Model Assignment
Input: remaining available memory size set $M$, remaining computational resource set $E$, the edge device set $D$, the sub-model set $Mo$
@@
To address the optimization problem expressed in Eq. 1, we propose a greedy search algorithm for assigning Vision Transformer sub-models to edge devices. As shown in Algorithm3, the sub-models are first sorted based on their energy consumption (computation overhead). ED-ViT assigns the most computation-intensive sub-model first based on their model sizes, which is proportional to the computation overhead as in Section 3. The algorithm then iteratively assigns the remaining sub-models to maximize the system’s available energy. Initially, the device with the highest computational power is selected. If the remaining memory and energy can accommodate the sub-model, we update the device’s available memory and energy. Otherwise, if the sub-model exceeds the device’s memory capacity, the memory-exhausted device is removed from the set. If no devices remain, it indicates that the current pruning results prevent deployment of all sub-models. In this case, the algorithm terminates, and the ED-ViT framework re-prunes the sub-models based on a new head pruning parameter, as described in Algorithm 1. Finally, the algorithm outputs the model assignment plan $MA$, representing the mapping of sub-models to edge devices.
As described in Section 3, the problem of Vision Transformer sub-model partitioning and assignment can be formulated as a 0-1 knapsack problem, where each edge device has varying available memory and energy. Each sub-model is responsible for a specific set of classes, and multiple sub-models can be deployed on a single device. We perform a collaborative optimization of partitioning the Vision Transformer model into multiple sub-models, as shown in Algorithm 1, and deploying these sub-models across edge devices using a greedy search assignment mechanism (Algorithm 3). This approach provides a relatively optimal solution to the formulated problem. Our extensive experiments demonstrate the effectiveness of our framework design and algorithms.
4.5. Model Fusion
In the result fusion phase, each sub-model on the edge devices processes inputs and extracts corresponding features. The server aggregates the generated features through concatenation and feeds them into an MLP to produce the final prediction. Notably, the MLP for result fusion requires training only once after all sub-models have been trained.
@@ 4.5. Model Fusion
5. Experiments
5.1. Experiments settings
Datesets. Considering the versatile applicability of the framework across various scenarios, we select three computer vision datasets (CIFAR-10 (Krizhevsky, 2009), MNIST (LeCun et al., 1998), and Caltech256 (Griffin et al., 2007)) and two audio recognition datasets (GTZAN (Tzanetakis and Cook, 2002) and Speech Command (Warden, 2018)) to construct the classification tasks for our experiments. We choose these datasets based on the real-world tasks as explained in Appendix A. For all the computer vision datasets, we resize the sample to $224\times 224\times 3$ to support various datasets and downstream tasks via a similar data structure without loss of generality; for the audio recognition datasets, we resize the sample to $224\times 224\times 1$ with the same aim. The details of the datasets and the preprocessing methods are shown in Appendix C.
@@ 5.1. Experiments settings
Figure 3. Performance metrics of Split ViT-Base models on CIFAR-10, MNIST, Caltech dataset. Note that (a) shows the accuracy results; (b) shows the latency results, the dotted lines represent the latency of the original ViT-Base model, and (c) shows the total memory sizes for all the sub-models. All the experiment results are collected on Raspberry Pi-4B.
Figure 4. Performance metrics of Split ViT-Base models on GTZAN and Speech Command dataset.
5.2. Experiments on Computer Vision Datasets
We evaluate our approach using CIFAR-10, MNIST, and Caltech image datasets. The original model size is 327.38 MB. Figure 3 shows the accuracy, inference latency, and memory usage of the ViT-Base model under the ED-ViT framework, with 1 to 10 edge devices. With only one edge device, we apply model compression by pruning Vision Transformer without decomposition. All experiments are conducted with a total memory budget of 180MB across devices, ensuring fair comparisons.
@@ 5.2. Experiments on Computer Vision Datasets
In terms of total memory usage, ED-ViT provides effective splitting and assignment strategies. As the number of edge devices increases, the memory size of each sub-model decreases, reducing computation overhead, and demonstrating that many complex model designs and computational operations are redundant for problem-solving. In the 10-edge device setting, the model size on the CIFAR-10 dataset is reduced to just 9.6MB, achieving a size reduction of up to 34.1 times compared to the original model.
@@
Figure 5. Performance metrics of Split ViT-Small and ViT-Large models on CIFAR-10, Caltech dataset.
5.3. Experiments on Audio Recognition Datasets
@@
We use the GTZAN and Speech Command audio datasets to evaluate the performance of our framework. The original model sizes of Vision Transformer for GTZAN and Speech Command is 325.88MB. Figure 4 presents the accuracy, inference latency, and total memory size of the ViT-Base model as implemented by the ED-ViT framework, similar to the experiments with the computer vision datasets. We still set the memory budget to 180MB.
@@
The results show that as the number of edge devices increases, ED-ViT is able to maintain the accuracy, delivering robust performance. For GTZAN, accuracies are consistently above 84%, and for the Speech Command dataset, accuracies are above 90%. Similar to the results on the computer vision datasets, the inference latency decreases as the number of edge devices increases. Notably, the latency for the original model is 32.16 seconds on the GTZAN dataset, which is 25.13 times the smallest latency (1.28s) and 3.37 times the highest latency (9.55s). This substantial latency reduction trend is consistent across both datasets. Regarding total memory size, all configurations remain within the set limits. As the number of edge devices increases, the memory size of each sub-model decreases, and the computation overhead is similarly reduced. In the 10-edge device setting, for each model on the GTZAN dataset, the size is reduced to only 9.35MB, achieving a reduction of up to 34.85 times compared to the original model. Similar results are observed for the Speech Command dataset.
5.4. Experiments on Different Vision Transformer Model Structures
@@
We also select two complex datasets (CIFAR-10 and Caltech) to test different Vision Transformer structures for low-power video analytics tasks. The original model sizes of ViT-Small and ViT-Large are 82.71MB, 1,157MB, respectively. Figure 5 presents the accuracy, inference latency, and total memory size of the ViT-Small and ViT-Large models as implemented by the ED-ViT framework, similar to Figure 3. We increase the total memory size limit for ViT-Large to 600MB and decrease the limit for ViT-Small to 50MB.
@@
The results show that as the number of devices increases, the accuracy remains relatively consistent, again showing robust performance. For ViT-Small, the accuracy is over 76.5% on the CIFAR-10 dataset and over 77.39% on Caltech across all settings; for ViT-Large, the accuracy is over 86% on the CIFAR-10 dataset and over 90.48% on Caltech in all settings. In most cases, the accuracy fluctuation for the final fusion prediction remains within a variance of less than one percentage point. The accuracy for ViT-Small is lower than that of ViT-Base, while ViT-Large achieves higher accuracy than ViT-Base, corresponding to the difference in parameter counts. Generally, the more parameters, the better the accuracy. As the number of edge devices increases, the latency decreases for both settings, similar to ViT-Base. The latency for ViT-Small is lower than that of ViT-Base, as ViT-Small requires less computational power, while the latency for ViT-Large is higher due to its larger size. In terms of memory size, in the 10-edge device setting, for each model on the CIFAR-10 dataset, the size for ViT-Small is 2.58MB, achieving a reduction of up to 32.06 times compared to the original model. Similarly, for ViT-Large, the size is 18.73MB, which also achieves a 61.77-fold reduction compared to the original model size. Note that for the ViT-Small on the CIFAR-10 and CalTech, the input size and the output size are the same; thus, their latency and total memory size on the edge devices are also the same. Similar results are observed across both datasets for ViT-Small and ViT-Large.
@@
Table 2. The accuracy results of splitting CNN and SNN versus ED-ViT with ViT-Base on CIFAR-10 dataset.
@@
5.5. Comparison with Baseline Methods: Splitting CNN and SNN
Vision Transformer achieves better accuracy compared to traditional CNN and SNN models. However, the performance of these models on edge devices has not been directly compared before. Nnfacet (Chen et al., 2021, 2023) proposes a method to split CNNs across multiple edge devices, employing a filter pruning technique (Hu et al., 2016), which differs from our approach. Yu et al. (2024) utilize the convolutional spiking neural network (CSNN) (Deng and Gu, 2021) to transform CNNs into SNNs, using a similar strategy. Both methods focus on VGGNet (Simonyan and Zisserman, 2014) networks and are channel-wise methods. In our experiments, the baseline model for these methods is VGGNet-16 in their papers, which also has a memory size similar to ViT-Base and achieves the best original results for comparison. We follow the hyper-parameters in their papers to get the results.
Figure 6. Performance metrics of splitting method with CNN, SNN, and ViT-Small models on CIFAR-10 dataset with 10 edge devices.
The accuracy results on the CIFAR-10 dataset are presented in Table 2. Based on the results, we observe that CNN outperforms SNN, while our ED-ViT method for ViT-Base yields better accuracy than both CNN and SNN approaches.
In addition to the accuracy results, we also compare inference latency and total memory size of the three methods when the number of edge devices is 10. These results are shown in Figure 6. Here, we selected ViT-Small, which has the lowest performance among the Vision Transformer structures. Based on the results, our ED-ViT method achieves the best accuracy compared to SNN and CNN, while its inference latency is much lower than SNN and CNN. Furthermore, the total memory size of ED-ViT is significantly lower than CNN and is comparable to SNN (as SNN is known for its small model size). This experiment demonstrates that deploying Vision Transformer onto edge devices can meet latency and memory constraints while delivering superior accuracy results.
Table 3. The impact of retraining for CIFAR-10 dataset on ViT-Base.

### v2

@@ IV Methodology
This section describes the design of the ED-ViT framework proposed to solve the optimization problem outlined in (1). We first explain the main workflow of ED-ViT and then provide detailed descriptions of the four key steps involved.
IV-A Design Overview
@@
As illustrated in Fig. 1, ED-ViT leverages the unique characteristics of Vision Transformer and the collaboration of multiple edge devices. The framework involves $N$ concurrent edge devices for distributed inference alongside a lightweight MLP aggregation to derive the final classification results. Initially, the original Vision Transformer is trained on the entire dataset to achieve high test accuracy for the classification task. The ED-ViT framework is composed of four main components: model splitting, pruning, assignment, and fusion. During model splitting, the Vision Transformer model is divided into sub-models, each responsible for a subset of classes. To reduce computation overhead, these sub-models are further pruned using model pruning techniques. Subsequently, the sub-models are assigned to the appropriate edge devices, taking the optimization problem into consideration. Finally, the aggregation device fuses the outputs from the edge devices to produce the final inference results. The specific details of each component are provided below.
@@
5:     $C_{i}$ is determined randomly.
6:  until $\forall C_{a},C_{b}\in C,\big{|}|C_{a}|-|C_{b}|\big{|}\leq 1$.
7:  while $flag_{r}$ is $True$ do
8:     for $i$ in $N$ do
@@
10:     end for
11:     $MA=\phi$
12:     if $\sum_{j}m_{i}\leq bu$ then
13:        $MA=greedySearchAssign(E,M,\text{Model}_{i},D)$.
14:     end if
15:     if  $MA\neq\phi$  then
16:        $flag_{r}=False$.
17:     else
18:        $hp_{j}=hp_{j}+1$ where Modelj has the biggest memory size.
19:     end if
20:  end while
21:  $concat_{out}$ = $concat$(Model1($\mathbf{X}$),…, ModelN($\mathbf{X}$)).
22:  $MLP$ = $train$($concat_{out}$, $\mathbf{y}$).
23:  return $MLP$, Model1, …, ModelN.
@@
IV-B Model Splitting
In the original Vision Transformer, different heads contribute to learning and inferring from the samples. However, for certain classes, maintaining all the connections between the heads can be redundant. As a result, ED-ViT prunes these connections and reconstructs the heads, with more retained heads leading to more parameters and connections being preserved. As illustrated in Algorithm 1, each Vision Transformer sub-model undergoes pruning based on a head number threshold and its associated categories, following a relatively equitable workload distribution. Subsequently, a greedy search mechanism is used to identify the most suitable edge device model assignment plan for deploying a particular sub-model, considering both energy and memory constraints. If the total memory size exceeds the budget or no suitable plan is found, an iterative approach is applied to adjust the number of heads for the sub-model with the biggest memory size to be pruned, repeating the allocation process until all sub-models are successfully assigned to edge devices. The pruning and the greedy assignment methods are shown as Algorithm 2 and Algorithm 3, located in Section IV-C and  IV-D, respectively.
IV-C Model Pruning
We believe that reducing the computational burden of Vision Transformer will significantly contribute to lowering inference latency in distributed edge device settings. We focus on the original ViT architecture [9], chosen for its simplicity and well-defined design space, focusing on redistributing the dimensionality across different blocks to achieve a more balanced tradeoff between computational efficiency and accuracy, as shown in Fig. 2.
Analysis of Prunable Parameters: The main prunable components in a ViT block, as illustrated in Fig. 2.
@@ IV-C Model Pruning
Pruning Process: As illustrated in Fig. 2, The pruning process is carried out in stages, with each stage focusing on one of the prunable components. We compute the KL-Divergence between the output distributions of the original model and the pruned model to evaluate the importance of each component, as follows:
@@
We focus on pruning the channels of the residual connections (as shown in red) in the first stage. Using KL-Divergence, we identify and prune the channels that contribute the least, reducing the dimensionality from $d$ to $s\times d$, and the pruning factor $s$ controls the degree of reduction in the parameters. We use $j$-th sub-model as an example: we set $s=(h-hp_{j})/h$, effectively controlling the extent of the pruning and parameter reduction. This helps to streamline the flow of information between layers without significantly affecting model performance. Then, instead of directly removing entire heads in the MHSA module, we prune the least important dimensions within the query, key, and value projections ($d_{q}$, $d_{k}$, and $d_{v}$) across multiple heads. This process effectively reduces the total number of heads to $s\times h$, without entirely discarding any head, thus maintaining a balanced representation of the attention mechanism while reducing its complexity. The dimensionality of the projections is scaled accordingly to reflect the merging and pruning process, ensuring that the model retains its ability to capture token interactions. The final stage involves pruning the hidden dimension $c$ in the FFN, as shown in blue. By calculating KL-Divergence, we identify the least important neurons and reduce the hidden dimension from $c$ to $s\times c$. Following each pruning stage, the model is fine-tuned to recover any performance loss that may result from the parameter reduction. This ensures that the pruned model achieves a similar level of accuracy as the original model while requiring fewer computational resources.
@@ IV-D Model Assignment
Input: remaining available memory size set $M$, remaining computational resource set $E$, the edge device set $D$, the sub-model set.
@@
To address the optimization problem expressed in (1), we propose a greedy search algorithm for assigning Vision Transformer sub-models to edge devices. As shown in Algorithm3, the sub-models are first sorted based on their energy consumption. ED-ViT assigns the most computation-intensive sub-model first based on their model sizes, which is proportional to the computation overhead as in Section III. The algorithm then iteratively assigns the remaining sub-models to maximize the system’s available energy. Initially, the device with the highest computational power is selected. If the remaining memory and energy can accommodate the sub-model, we update the device’s available memory and energy. Otherwise, if the sub-model exceeds the device’s memory capacity, the memory-exhausted device is removed from the set. If no devices remain, it indicates that the current pruning results prevent deployment of all sub-models. In this case, the algorithm terminates, and the ED-ViT framework re-prunes the sub-models based on a new head pruning parameter, as described in Algorithm 1. Finally, the algorithm outputs the model assignment plan $MA$, representing the mapping of sub-models to edge devices.
As described in Section III, the problem of Vision Transformer sub-model partitioning and assignment can be formulated as a 0-1 knapsack problem, where each edge device has varying available memory and energy. Each sub-model is responsible for a specific set of classes, and multiple sub-models can be deployed on a single device. We perform a collaborative optimization of partitioning the Vision Transformer model into multiple sub-models, as shown in Algorithm 1, and deploying these sub-models across edge devices using a greedy search assignment mechanism , as shown in Algorithm 3. This approach provides a relatively optimal solution to the formulated problem. Our extensive experiments demonstrate the effectiveness of our framework design and algorithms.
IV-E Model Fusion
In the result fusion phase, each sub-model on the edge devices processes inputs and extracts corresponding features. The aggregation edge device aggregates the generated features through concatenation and feeds them into an MLP to produce the final prediction. Notably, the MLP for result fusion requires training only once after all sub-models have been trained.
@@ IV-E Model Fusion
V Experiments
V-A Experiments settings
Datasets. Considering the versatile applicability of the framework across various scenarios, we select three computer vision datasets (i.e., CIFAR-10 [55], MNIST [56], and Caltech256 [57]) and two audio recognition datasets (i.e., GTZAN [58] and Speech Command [59]) to construct the classification tasks for our experiments. For all the computer vision datasets, we resize the sample to $224\times 224\times 3$ to support various datasets and downstream tasks via a similar data structure without loss of generality; for the audio recognition datasets, we resize the sample to $224\times 224\times 1$ with the same aim.
Figure 3: Our 5-device example experimental prototype utilizes a switch and Raspberry Pi 4B devices, with one dedicated to the fusion model and the other four allocated to sub-models.
@@ V-A Experiments settings
Figure 4: Performance metrics of Split ViT-Base models on CIFAR-10, MNIST, Caltech dataset. Note that (a) shows the accuracy results; (b) shows the latency results, the dotted lines represent the latency of the original ViT-Base model, and (c) shows the total memory sizes for all the sub-models. All the experiment results are collected on Raspberry Pi-4B.
Figure 5: Performance metrics of Split ViT-Base models on GTZAN and Speech Command dataset.
V-B Experiments on Computer Vision Datasets
We evaluate our approach using CIFAR-10, MNIST, and Caltech image datasets. The original model size is 327.38 MB. Fig. 4 shows the accuracy, inference latency, and memory usage of the ViT-Base model under the ED-ViT framework, with 1 to 10 edge devices. With only one edge device, we apply model compression by pruning Vision Transformer without decomposition. All experiments are conducted with a total memory budget of 180MB across devices, ensuring fair comparisons.
@@ V-B Experiments on Computer Vision Datasets
In terms of total memory usage, ED-ViT provides effective splitting and assignment strategies. Note that as the number of retained heads increases, the memory size grows quadratically. For one edge device, retaining more heads could exceed the budget. However, in a two-device setting, each sub-model retains a similar number of heads, ensuring the total memory usage remains within the budget. This explains the spike in total memory sizes with two edge devices, as shown in Fig. 4. As the number of edge devices increases from 3 to 10, the memory size of each sub-model decreases, reducing computation overhead and demonstrating that many complex model designs and computational operations are redundant for problem-solving. In the 10-edge device setting, the model size on the CIFAR-10 dataset is reduced to just 9.60MB, achieving a size reduction of up to 34.1 times, compared to the original model from ED-ViT.
@@
Figure 6: Performance metrics of Split ViT-Small and ViT-Large models on CIFAR-10, Caltech dataset.
V-C Experiments on Audio Recognition Datasets
@@
We use the GTZAN and Speech Command audio datasets to evaluate the performance of our framework. The original model sizes of Vision Transformer for GTZAN and Speech Command is 325.88MB. Fig. 5 presents the accuracy, inference latency, and total memory size of the ViT-Base model as implemented by the ED-ViT framework, similar to the experiments with the computer vision datasets. We still set the memory budget to 180MB.
@@
The results show that as the number of edge devices increases, ED-ViT is able to maintain the accuracy, delivering robust performance. For GTZAN, accuracies are consistently above 84%, and for the Speech Command dataset, accuracies are above 90%. Similar to the results on the computer vision datasets, the inference latency decreases as the number of edge devices increases. Notably, the latency for the original model is 32.16 seconds on the GTZAN dataset, which is 25.13 times the smallest latency (1.28s) and 3.37 times the highest latency (9.55s). This substantial latency reduction trend is consistent across both datasets. Regarding total memory size, all configurations remain within the set limits. As the number of edge devices increases, the memory size of each sub-model decreases, and the computation overhead is similarly reduced. In the 10-edge devices setting, for each model in the GTZAN dataset, the size is reduced to only 9.35MB, achieving a reduction of up to 34.85 times compared to the original model. Similar results are also observed in the Speech Command dataset from ED-ViT.
TABLE II: The FLOPs for sub-models on different datasets when using ViT-Base.
Dataset The Number of Edge Devices
Original 2 3 5 10
CIFAR-10 16.86G 4.25G 1.90G 1.08G 0.48G
GTZAN 16.79G 4.20G 1.88G 1.059G 0.46G
V-D Overhead of Computation and Communication
We use FLOPs to simulate the energy consumption on the edge devices. Table II presents the FLOPs for each edge device on CIFAR (computer vision/video) and GTZAN (audio recognition) datasets with different numbers of edge devices using the ViT-Base model. The FLOPs for the original model on the CIFAR-10 and GTZAN datasets are 16.86G and 16.79G, respectively. As the number of edge devices increases, the FLOPs decrease across all datasets, and the experimental results are consistent with the parameter counts. These findings demonstrate that ED-ViT significantly reduces computation overhead and saves energy for the edge devices.
When the number of edge devices increases from 1 to 10 using ViT-Base across all the datasets, the size of features for communication on each sub-model decreases from 1536 bytes to 512 bytes. Compared with the original image size (150528 bytes), our method could greatly reduce the communication overhead to 294 times. The maximal communication time for one edge device is 5.86ms, which is acceptable in the practical situation. The results also show that the inferences on sub-models and the fusion model take up most of the latency (order of seconds) in Section V-B and Section V-C.
V-E Experiments on Different Model Structures
@@
We also select two complex datasets (e.g, CIFAR-10 and Caltech) to test different Vision Transformer structures for low-power video analytics tasks. The original model sizes of ViT-Small and ViT-Large are 82.71MB, 1,157MB, respectively. Fig. 6 presents the accuracy, inference latency, and total memory size of the ViT-Small and ViT-Large models as implemented by the ED-ViT framework, similar to Fig. 4. We increase the total memory size limit for ViT-Large to 600MB and decrease the limit for ViT-Small to 50MB.
@@
The results show that as the number of devices increases, the accuracy remains relatively consistent, again showing robust performance. For ViT-Small, the accuracy is over 76.5% on the CIFAR-10 dataset and over 77.39% on Caltech across all settings; for ViT-Large, the accuracy is over 86% on the CIFAR-10 dataset and over 90.48% on Caltech in all settings. In most cases, the accuracy fluctuation for the final fusion prediction remains within a variance of less than one percentage point. The accuracy for ViT-Small is lower than that of ViT-Base, while ViT-Large achieves higher accuracy than ViT-Base, corresponding to the difference in parameter counts. Generally, the more parameters, the better the accuracy. As the number of edge devices increases, the latency decreases for both settings, similar to ViT-Base. The latency for ViT-Small is lower than that of ViT-Base, as ViT-Small requires less computational power, while the latency for ViT-Large is higher due to its larger size. In terms of memory size, in the 10-edge device setting, for each model on the CIFAR-10 dataset, the size for ViT-Small is 2.58MB, achieving a reduction of up to 32.06 times compared to the original model. Similarly, for ViT-Large, the size is 18.73MB, which also achieves a 61.77-fold reduction compared to the original model size.
Note that for the ViT-Small on the CIFAR-10 and CalTech, the input size and the output size are the same; thus, their latency and total memory size on the edge devices are also the same. Similar results are observed across both datasets for ViT-Small and ViT-Large.
@@
TABLE III: The accuracy results of splitting CNN and SNN versus ED-ViT with ViT-Base on CIFAR-10 dataset.
@@
V-F Comparison with Baseline Methods: Split-CNN and Split-SNN
Vision Transformer achieves better accuracy compared to traditional CNN and SNN models. However, the performance of these models on edge devices has not been directly compared before. Nnfacet [2] proposes a method to split CNNs across multiple edge devices, employing a filter pruning technique [65], which differs from our approach. EC-SNN [3] utilize the convolutional spiking neural network (CSNN) [7] to transform CNNs into SNNs, using a similar strategy. Both methods focus on VGGNet [5] backbone networks and are channel-wise methods. In our experiments, the baseline model for these methods is VGGNet-16 in their papers, which also has a memory size similar to ViT-Base and achieves the best original results for comparison. We follow the hyper-parameters in their papers to conduct the experiments.
Figure 7: Performance of splitting method with CNN, SNN, and ViT-Base models on CIFAR-10 dataset with 10 edge devices.
The accuracy results on the CIFAR-10 dataset are presented in Table III. Based on the results, we observe that CNN outperforms SNN, while our ED-ViT method for ViT-Base yields better accuracy than both CNN and SNN approaches. The original accuracies of ViT-Base, CNN, and SNN are 98.12%, 93.64%, and 93.56%, respectively. Their average accuracy losses are 11.16%, 8.5% and 10.15% across various device numbers. Due to the inherently high accuracy of Vision Transformer, we admit that its accuracy drop is slightly higher than CNN and SNN. This is precisely why we leverage Vision Transformer: to employ its exceptional performance. Our method maintains a comparable accuracy drop with a 28.9× size reduction, achieving up to 4.06% and 5.55% higher accuracy than state-of-the-art CNN-based and SNN-based methods shown in Table III.
In addition to the accuracy results, we also compare inference latency and total memory size of the three methods when the number of edge devices is 10. These results are shown in Fig. 7. Based on the results, our ED-ViT method achieves the best accuracy compared to SNN and CNN, while its inference latency is much lower than SNN (4.36 times) and CNN (2.70 times). Furthermore, the total memory size of ED-ViT is significantly lower than CNN and is comparable to SNN, since SNN is known for its small model size. This experiment demonstrates that deploying Vision Transformer onto edge devices can meet latency and memory constraints while delivering superior accuracy results.
TABLE IV: The impact of retraining for CIFAR-10 dataset on ViT-Base of ED-ViT.